OpenStudy (anonymous):

Limit[(Cos[t + h] - Cos[t])/h, h -> 0]

OpenStudy (phi):

one way to do this is first use the identify cos( a+b) = cos(a) cos(b) - sin(a) sin(b) to write cos(t+h) = cos(t) cos(h) - sin(t) sin(h) next, we can use the series definition of cos and sin see https://en.wikipedia.org/wiki/Trigonometric_functions#Series_definitions in particular, write cos(h) = 1 - h^2/2 ..... and higher order terms sin(h)= h - h^3/6 .... and higher order terms use those definitions in \[ \frac{\cos(t) \cos(h) - \sin(t) \sin(h) - \cos(t)}{h} \] you should be able to divide h evenly into the numerator then take the limit as h->0

OpenStudy (anonymous):



OpenStudy (phi):

replace cos(h) with ( 1 - h^2/2 +....) and sin(h) with ( h - h^3/6 ....)

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

how about \[\lim_{h \rightarrow 0}\frac{ \cos h - 1 }{ h }=0\text{, and }\lim_{h \rightarrow 0}\frac{ \sin h }{ h }=1\] that is hopefully something they've seen previously

OpenStudy (anonymous):

so the answer will be -sin(t)

OpenStudy (anonymous):

yep

OpenStudy (jdoe0001):

pgpilot326 I sorta went that route and... thus far have... \(\bf lim_{h \to 0}\cfrac{cos(t+h)-cos(t)}{h}\\ \quad \\ \quad \\ \cfrac{cos(t+h)-cos(t)}{h}\implies \cfrac{cos(t)cos(h)-sin(t)sin(h)-cos(t)}{h}\\ \quad \\ \textit{getting common factor atop}\\ \quad \\ \cfrac{cos(t)[cos(h)-1]-sin(t)sin(h)}{h}\\ \quad \\ \cfrac{cos(t)[cos(h)-1]}{h} - \cfrac{sin(t)sin(h)}{h}\\ \quad \\ \cfrac{cos(t)}{h}\cdot \cfrac{cos(h)-1}{h}- \cfrac{sin(t)}{h}\cdot \cfrac{sin(h)}{h} \implies -\cfrac{sin(t)}{h}\)

OpenStudy (anonymous):

no that last line is incorrect. shouldn't have h twice in the denominator

OpenStudy (jdoe0001):

ohh yeah!

OpenStudy (anonymous):

oh using l hopital's rule right?

OpenStudy (jdoe0001):

man... yes indeed

OpenStudy (anonymous):

sweet thanks guys... =)

OpenStudy (jdoe0001):

\(\bf lim_{h \to 0}\cfrac{cos(t+h)-cos(t)}{h}\\ \quad \\ \quad \\ \cfrac{cos(t+h)-cos(t)}{h}\implies \cfrac{cos(t)cos(h)-sin(t)sin(h)-cos(t)}{h}\\ \quad \\ \textit{getting common factor atop}\\ \quad \\ \cfrac{cos(t)[cos(h)-1]-sin(t)sin(h)}{h}\\ \quad \\ \cfrac{cos(t)[cos(h)-1]}{h} - \cfrac{sin(t)sin(h)}{h}\\ \quad \\ \cfrac{cos(t)}{h}\cdot \cfrac{cos(h)-1}{h}- sin(t) \cdot \cfrac{sin(h)}{h} \implies -sin(t)\)

OpenStudy (anonymous):

1st term too

OpenStudy (jdoe0001):

hehe, you're right...

OpenStudy (jdoe0001):

\(\bf cos(t)\cdot \cfrac{cos(h)-1}{h}- sin(t) \cdot \cfrac{sin(h)}{h} \implies -sin(t)\)

OpenStudy (anonymous):

that's it, sans the limit

OpenStudy (phi):

Limit[(Cos[t + h] - Cos[t])/h, h -> 0] is going back to the definition of a derivative. \[ \frac{ \cos(t) \cos(h) - \sin(t) \sin(h) - \cos(t)}{h} \\ \frac{ \cos(t)( 1 - \frac{h^2}{2}+...) - \sin(t)(h - \frac{h^3}{3!}+...) - \cos(t)}{h} \\ \frac{ \cancel{\cos(t)} - \frac{h^2}{2} \cos(t)+... - h\sin(t) - \frac{h^3}{3!}\sin(t)+... - \cancel{\cos(t)} }{h} \\ - \frac{h}{2}\cos(t)+... - \sin(t) - \frac{h^2}{6} \sin(t)... \] now take the limit as h->0 and all the terms drop out except -sin(t)

OpenStudy (anonymous):

yes but i3lue probably hasn't seen series yet. how does i3lue know that the series representations of sin and cos are valid?

OpenStudy (anonymous):

are those factorial?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i see what pi is trying to say .... hahaha its been awhile since i did series lol =P

OpenStudy (anonymous):

yeah, what he's saying is perfectly valid

OpenStudy (agent0smith):

Limit[(Cos[t + h] - Cos[t])/h, h -> 0] why not apply l'hopitals rule from the start, if you're going to use it later? differentiate w.r.t. h...

OpenStudy (agent0smith):

That might not be perfectly valid, but it works... treating t as a constant. Limit[(Cos[t + h] - Cos[t])/h, h -> 0] becomes Limit[-sin(t+h))/1, h -> 0]

OpenStudy (agent0smith):

Now plug in h=0, and you get -sin(t)

OpenStudy (agent0smith):

I mean, t is a constant... it's h that's approaching zero, t is constant.

OpenStudy (anonymous):

Thanks =)

OpenStudy (agent0smith):

I mean i guess my way kinda circumvents the actual definition of the derivative, since you need to use actual derivative rules...